# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ IV STALC Technical Approach
Figure 2: STALC: Hierarchical planning architecture for a multi-robot team to traverse an environment. The high-level planner segments the environment into a topological graph structure and generates a plan on the graph for the robot team through the environment. The graph construction is informed by the mid-level planner which generates A paths through the environment. The mid-level planner then takes the output of the graph planner and assigns routes to each robot through the environment. Each group of robots moving together are then assigned roles in their sub-team (i.e., leader or follower). This information is sent to each robot’s low-level planner to generate and execute dynamically feasible, collision-free paths.
Our technical approach consists of a three-level planning hierarchy including a top-level graph-based planner capable of coordinating multi-robot teams while traversing a hazardous environment. The edges in our graph represent minimally visible paths between the graph’s nodes, which we generate with our mid-level planner using a visibility map and A search. At the lowest-level, we utilize Model Path Predictive Integral Control (MPPI) [33] to generate local plans, avoid obstacles, and follow the mid-range plan. All three levels share visibility and terrain maps to maintain consistent planning across the hierarchy.
@@
Fig. 2 shows a flow diagram of our hierarchical control approach. First, terrain information is used to compute visibility scores, relative to an expected observer distribution, across an area of interest. We use this visibility map to segment the environment into regions of cover (or low visibility), which we use as nodes in a graph structure. We then calculate minimum visibility A paths between the regions to serve as the edges in the graph. Between each node and edge pair, we compute overwatch scores based on the visibility to the edge from within the node. Together, this information forms a dynamic topological graph for the environment, where the edge weights vary with the state of the robot team. We encode this graph in our MIP graph planner to compute an optimal plan for the multi-robot team. To form our mid-level plan, we assign each robot a route, which corresponds to an A path, and a role in the team (i.e., leader or follower). Each robot then controls to the A paths using MPPI in our low-level planner, performs overwatch, and/or moves in formation with their team members. The following sections will outline the algorithms that comprise each block of our overall hierarchical planner.
V Graph Generation
V-A Visibility Map
To enable low-visibility planning for the multi-robot team, we construct a visibility map. We assume prior knowledge of a Digital Elevation Model (DEM) of the environment and a probability distribution, $\mathcal{O}$, over $\mathbb{R}^{3}$ representing the expected observer position. We do not make assumptions about the form of $\mathcal{O}$ except that we are able to sample from it. In the experimental evaluation that follows, we utilize Gaussian distributions (i.e., $\mathcal{O}\sim N(\mu,\Sigma)$).
For a given DEM and observer distribution, we segment the environment into regions of low and high visibility from the perspective of the observer. We accomplish this by adapting the notion of a viewshed[34] to the context of a distribution over observer positions. The viewshed of an area is a binary mask marking the regions of a DEM that are visible from a given observer position. Our goal is to estimate the probability that each point in the environment can be seen by the observer.
[…]
The calculation outlined above does not consider the distance from the observer when computing the probability of detection; however, we expect visibility to decrease with increasing distance. To model this effect, we use the distance transform to compute the Euclidean distance from each cell in the map to the nearest observer distribution. For distributions with infinite support, such as a normal distribution, we must choose a “boundary” for the distribution in order to compute the distance transform. In this work, we use the 2-$\sigma$ ellipse. We compute a weight for each cell in the map, such that the probability of detection is 0 beyond the max range and approaches the visibility map value for locations closer to an observer. Thus, we define our visibility map, $\mathcal{P}_{vis}(x,y)$, as follows.
[…]
In this calculation, $d(x,y)$ is the distance to the nearest observer and $d_{max}$ is the maximum sight range of the observer.
Algorithm 1 Topological Graph Construction
0:
$\mathcal{D}$ Digital Elevation Model
$\mathcal{C}$ Obstacle map
$\mathcal{O}$ Observer distribution
$N$ Number of samples from observer distribution
$\nu$ Visibility threshold
$\Xi_{min}$ Minimum cover region size
$\Xi_{max}$ Maximum cover region size
$\lambda_{p}$ Visibility cost weight
0:
$V$ Graph nodes (cover region centroids)
$E$ Graph edges
$W_{vis}$ Visibility weight adjacency matrix
$W_{ow,V}$ Overwatch weight matrices for each node $v\in V$
1:  $\mathcal{P}_{vis}\leftarrow$ ComputeVisibilityMap($\mathcal{D}$, $\mathcal{O}$, $N$)
2:  $cover\leftarrow$ ThresholdVisibility($\mathcal{P}_{vis}$, $\nu$)
3:  $cover\leftarrow$ RemoveObstacles($\mathcal{C}$, $𝑜𝑣𝑒𝑟c$)
[…]
5:  $regions\leftarrow$ SplitRegions($regions$, $\Xi_{max}$)
6:  $V\leftarrow$ ComputeCentroids($\mathcal{C}$, $regions$)
7:  $paths\leftarrow$ ComputePaths($\mathcal{P}_{vis}$, $\mathcal{C}$, $regions$, $V$, $\lambda_{p}$)
8:  $\mathcal{N}_{vis}\leftarrow-\log(\max(1-\mathcal{P}_{vis},\epsilon))$
9:  $n_{V}\leftarrow$ Size($regions$)
10:  $W_{vis}\leftarrow\textbf{0}_{n_{V}\times n_{V}}$
11:  $E\leftarrow\emptyset$
12:  for all $\{(s,g):path\}\in paths$ do
13:     $W_{vis}[s,g]\leftarrow$ PathCost($\mathcal{N}_{vis}$, $path$)
14:     $E\leftarrow E\cup(s,g)$
15:  end for
16:  for $v\in V$ do
17:     $\mathcal{O}_{ow,v}\leftarrow$ UniformDistribution($regions[v]$)
18:     $\mathcal{P}_{ow,v}\leftarrow$ ComputeVisibilityMap($\mathcal{D}$, $\mathcal{O}_{ow,v}$, $N$)
19:     $\mathcal{N}_{ow,v}\leftarrow-\log(\max(1-\mathcal{P}_{ow,v},\epsilon))$
20:     $W_{ow,v}\leftarrow\textbf{0}_{n_{V}\times n_{V}}$
21:     for all $\{(s,g):path\}\in paths$ do
22:        $W_{ow,v}[s,g]\leftarrow$ PathCost($\mathcal{N}_{ow,v}$, $path$)
23:     end for
24:  end for
25:  return  $V,E,W_{vis},W_{ow,V}$
Algorithm 2 ComputePaths
0:
$\mathcal{P}_{vis}$ Visibility map
$\mathcal{C}$ Obstacle map
$regions$ Cover regions
$V$ Graph nodes
$\lambda_{p}$ Visibility cost weight
0:
$paths$ Map: $(start,goal)\mapsto path$
1:  $paths\leftarrow$ NewMap()
2:  $is\_start\leftarrow\{v:\textbf{false}{}\;\forall v\in V\}$
3:  $is\_goal\leftarrow\{v:\textbf{false}{}\;\forall v\in V\}$
4:  $path\_to\leftarrow\{v:\textbf{null}\;\forall v\in V\}$
5:  $path\_from\leftarrow\{v:\textbf{null}\;\forall v\in V\}$
6:  for all $s\in V$ do
7:     for all $g\in V$ do
8:        if $s=g$ then
9:           continue
10:        end if
11:        $path\leftarrow$ ComputeSinglePath($\mathcal{C}$, $\mathcal{P}_{vis}$, $s$, $g$, $\lambda_{p}$)
12:        if  PathIsRedundant($path$, $s$, $g$)  then
13:           $path\_to[g]\leftarrow$ MinCost($path$, $path\_to[g]$)
14:           $path\_from[s]\leftarrow$ MinCost($path$, $path\_from[s]$)
15:        else
16:           $paths[\{s,g\}]\leftarrow path$
17:           $is\_start[s]\leftarrow$ true
18:           $is\_goal[g]\leftarrow$ true
19:        end if
20:     end for
21:  end for
22:  $paths\leftarrow$ ReconnectNodes($is\_start$, $is\_goal$,             $path\_from$, $path\_to$)
23:  return  $paths$
V-B Constructing a Topological Graph
In order to construct a topological graph for planning, we use the visibility map to compute a set of nodes, $V$, and edges, $E$, together with a score for the visibility of each edge, represented by a weighted adjacency matrix, $W_{vis}$. We also compute an overwatch matrix for each node, $W_{ow,v}$ for $v\in V$, which is a weighted adjacency matrix where element $W_{ow,v}(j,k)$ denotes the overwatch that a team of robots positioned at node $v$ can provide for a team traversing edge $(j,k)$. In Algorithm 1, we present our algorithm to construct the components of the topological graph. The following subsections describe the key steps (represented in the algorithm as functions) in more detail.
V-B1 ComputeVisibilityMap
The visibility map computation uses an efficient viewshed algorithm [34] to generate $N$ viewshed maps sampled from the observer distribution $\mathcal{O}$. We then use equations (1) and (2) to compute the probability of detection for each cell in the map. These calculations form our visibility map $\mathcal{P}_{vis}$.
V-B2 ThresholdVisibility
To identify the regions of low visibility (which we call cover regions) from the visibility map, we threshold the map to extract regions below a pre-defined visibility cutoff, $\nu$.
V-B3 RemoveObstacles
After thresholding, we additionally remove any regions of the map that are known a priori to be non-traversable.
V-B4 ConnectedComponents
We then perform a connected-components search to find all the discrete regions whose sizes exceed a threshold value, $\Xi_{min}$.
Figure 4: A visibility map overlaid on satellite imagery of the environment: cover regions are outlined in red and node numbers are located within the regions near the centroid of each region.
Figure 5: Pruned A paths between regions of cover.
V-B5 SplitRegions
For some maps and observer positions, the cover regions generated using our method can be large and quite non-convex. Consequently, we enforce a maximum size, $\Xi_{max}$, for each cover region by iteratively finding the minimum-length cut that divides one region into two smaller ones, where one region has the maximum size [35]. These cuts are repeated until no region is larger than the maximum size.
V-B6 ComputeCentroids
The nodes of the topological graph are initially placed at the centroid of each cover region. We then shift them to ensure that the nodes lie inside the cover region (as the centroid could be outside of a non-convex region). Additionally, we ensure the node centers do not collide with any known obstacles. In Fig. 4, we overlay outlines of the cover regions with the visibility map on satellite imagery of the environment.
V-B7 ComputePaths
We define the algorithm for computing paths between cover regions in Algorithm 2. In particular, we generate a mapping $paths$ that stores the path between any two start and goal nodes. We initialize this mapping with the NewMap function.
V-B8 ComputeSinglePath
To identify redundant edges and compute the edge weights for the graph, we use an optimal path planning algorithm to compute a path between each pair of nodes in the graph. In this work, we use a standard A search where the heuristic cost, $h(b)$, is the Euclidean distance from the current position to the goal, $g(b)$ denotes the cost to follow the optimal path from the start to $b$, and the successor cost, $c(b,b^{\prime})$, is the cost to go from a cell in the map to a neighbor, with terms incorporating both the distance and the detection probability.
$c(b,b^{\prime})=||b-b^{\prime}||(1+\lambda_{p}\mathcal{N}_{vis}(b^{\prime}))$ (3)
$g(b^{\prime})=g(b)+c(b,b^{\prime})$ (4)
In (3), $\lambda_{p}$ is a scaling factor to weight the contribution of the visibility cost. Since the true cost-to-go is equal to the heuristic cost plus a non-negative visibility term, the heuristic $h(b)$ is trivially admissible.
V-B9 PathIsRedundant
When computing the set of graph edges and the cost to traverse each edge, we initially consider the topological graph to be fully-connected. In order to reduce the number of edges considered in the graph planner (Sec. VI), we prune the graph by eliminating redundant edges. A redundant edge is any edge connecting node $v_{i}$ to node $v_{j}$ where the optimal path from $v_{i}$ to $v_{j}$ passes through a cover region associated with another node $v_{k}\notin\{v_{i},v_{j}\}$. Whenever a redundant edge is removed, we compare the cost of the corresponding path with both the current minimum cost redundant path starting at $v_{i}$ and the current minimum cost redundant path ending at $v_{j}$ using the function MinCost and save the lower cost paths for later use in ReconnectNodes.
V-B10 ReconnectNodes
In rare circumstances, the pruning step leaves a node without any connections to the graph. In this case, we restore the lowest cost edge for which the disconnected node is a start node and the lowest cost edge where it is a goal node. An example of the pruned A paths is shown in Fig. 5.
V-B11 PathCost
When traversing a path, $B=\{b_{1},b_{2},...,b_{n}\}$, $b_{i}=[x_{i},y_{i}]^{T}$, through the environment, we consider the total probability of non-detection, $P_{nd}$, as the product of the probabilities of non-detection at each cell (assuming independence). This can be calculated for a visibility map $\mathcal{P}_{vis}$ or overwatch map $\mathcal{P}_{ow,v}$, so we use a general $\mathcal{P}$ in our equations.
$P_{nd}=\prod_{b\in B}(1-\mathcal{P}(b))$ (5)
Taking the negative $\log$ of this quantity, we find that we can compute a negative $\log$ non-detection probability by summing the contributions of each cell along the path.
$-\log(P_{nd})=\sum_{b\in B}-\log(1-\mathcal{P}(b))$ (6)
As the probability of detection approaches one, the cost to traverse a cell approaches infinity. In order to avoid infinite values in the optimization problem, we bound the maximum probability of detection at each cell to $1-\epsilon$ (for a small positive $\epsilon$). We call the map containing the negative $\log$ non-detection probabilities at each cell $\mathcal{N}$, or $\mathcal{N}_{vis}$ for visibility and $\mathcal{N}_{ow,v}$ for overwatch.
In the final step of the path cost calculation, we compute the weight for each edge $e=(v_{1},v_{2})$ that remained after the pruning step by considering the optimal path $B_{e}$.
$W(v_{1},v_{2})=\sum_{b\in B_{e}}\mathcal{N}(b)$ (7)
V-B12 Overwatch Opportunities
An overwatch opportunity represents the ability of a subset of robots in a cover region to provide support to another team as they traverse an edge in the graph. The process of computing an overwatch weight matrix, $W_{ow,v}$, which represents the benefit of performing overwatch, is very similar to the visibility map calculation; however, instead of sampling from the observer distribution, we sample uniformly from the cover region associated with the node $v$. We sample uniformly since we have hard boundaries for the cover region, whereas in the case of the observer, we may not know the boundaries, so we considered a Gaussian distribution. We call the uniform distribution associated with the node $v$ cover region $\mathcal{O}_{ow,v}$. This distribution is formed with the function UniformDistribution. Using this distribution, we produce a visibility map for an observer located at node $v$, which we call an overwatch map, $\mathcal{P}_{ow,v}$. We then apply equations (6) and (7) to $\mathcal{P}_{ow,v}$ to compute the overwatch score for each edge in the graph. We perform this calculation for each node resulting in a series of overwatch weight matrices $W_{ow,V}$ to be used in our topological graph. An example overwatch map is shown in Fig. 6.
V-C Topological Graph Refinement
We graphically depict the information embedded in our topological graph through the schematic representation shown in Fig. 7. The graph includes nodes for each cover region, edges representing the pruned A paths, and overwatch opportunities from the overwatch maps, denoted by arrows from the overwatch node to the edge that can be monitored.
When transforming the environment segmentation into this schematic representation, we use heuristics to refine the graph based on the robots’ capabilities and the priorities of the scenario. In particular, we set a maximum distance that can be traversed and remove edges from $E$ that do not meet this threshold. When necessary due to a particular scenario, we add vulnerable edges to nodes that became isolated from the maximum distance threshold by adding back the lowest-cost edge to $E$.
We add overwatch opportunities from $W_{ow,V}$ to our graph when a robot from a node can provide support to the robot(s) traversing an edge. To determine which overwatch opportunities are valuable to the planning problem, we enforce overwatch heuristics based on the scenario. First, we impose a maximum overwatch distance (e.g., due to sensor constraints) by checking the distance to each end point of an edge. We then add a scale factor to the overwatch weights based on the priorities of the scenario (e.g., if overwatch is more or less valuable). We consider providing overwatch to offer a cost reduction to traverse an edge, so we bound the ratio between the overwatch reduction and edge weight. Typically, we set these bounds to saturate at a 90% cost reduction due to overwatch (such that traversing an edge is never without cost) and we only consider the overwatch opportunity if it provides minimally a 40% cost reduction (based on what is considered worthwhile in the scenario). These thresholds and the scale factor on overwatch would be set based on the priorities of the scenario and we demonstrate different values for these parameters in our analysis.
VI Graph Planner
To solve for an optimal plan for a team of robots to navigate an environment while optimizing for our scenario objectives, we developed a MIP approach to capture the relationships in our topological graph. In particular, we consider our graph to be a dynamic topological graph, as it explicitly embeds the dynamic state of the robot team in the edge costs in the formulation of our optimization problem.
VI-A Mixed-Integer Programming Approach
In order to encode the dynamic topological graph structure in our MIP problem, we define decision variables that track the number of robots at each location (node or edge) in the graph. We then constrain the robots’ movement to the graph structure by considering the flow of robots between nodes and edges. We consider graph $G=(V,E,W_{E})$ with nodes $v\in V$, edges $e\in E$, and edge costs $w_{e}\in W_{E}:E\times X\rightarrow\mathbb{R}_{>0}$. The edge costs $w_{e}$ depend on the costs from the visibility map ($W_{vis}$), overwatch opportunities ($W_{ow,V}$), and the state of the robot team on the graph $x\in X$, where $X=\{p_{e_{1}},p_{e_{2}},\dots,p_{e_{nE}}\}$ and $p_{e_{j}}\in\mathbb{Z}_{>0}$ is the number of robots on edge $e_{j}$. We consider a fixed time horizon $t\in[1,n_{T}]$. Our overall objective is to position a team of robots at a set of target locations within this time horizon while minimizing time and visibility due to traversing by using our constructs defined in Sec. III. Thus, we define cost functions for the cost of traversing $C_{W_{e,t}}$ (based on the edge weights on the graph and reductions due to moving as a team), the cost reduction from overwatch $C_{\Omega_{{o},t}}$, and a cost for time $C_{T_{t}}$. Our cost functions depend on the positions of the robots relative to their teammates and reflect the dynamic element of the topological graphs. Specifically, the overall cost to traverse the graph decreases when the conditions for overwatch, moving in formation, or other general teaming are met.
We express a general form of our problem as follows, where we define an indicator function $\chi$, which checks that the state of the robot team is constrained to the graph structure, and starts and ends at positions defined by the sets $S$ and $D$.
$\min~{}$ $\sum_{t=1}^{n_{T}}\bigg{(}C_{T_{t}}+\sum_{e\in E}C_{W_{e,t}}+\sum_{{o}\in{O}}C_{\Omega_{{o},t}}\bigg{)}$ (8)
s.t. $\chi(x,G,S,D)=1$ (9)
In the following sections we expand on this formulation, defining specific cost functions and constraints. We consider a reconnaissance test scenario, however the algorithmic developments presented in this paper could be applied to a wide variety of scenarios.
A key innovation of this method is our ability to remove the standard dependence on the number of robots. Instead of tracking each robot’s path through the graph, we track the total number of robots at each node in the graph at each time step. This formulation abstracts away the assignment of paths for individual robots to a rapid post-processing step, ultimately significantly improving computation times. Furthermore, this formulation is flexible to a varying number of robots.
We define the parameters that we use to formulate the multi-robot MIP in Table I and the decision variables in Table II. Furthermore, we specify the constraints for the lower bounds (LB) and upper bounds (UB) of our decision variables in Table II.

### v2

@@ IV STALC Technical Approach
Figure 2: STALC: Hierarchical planning architecture for a multi-robot team to autonomously coordinate while traversing an environment. The high-level planner segments the environment into a topological graph structure and generates a plan on the graph for the robot team through the environment. The graph construction is informed by the mid-level planner which generates A paths through the environment. The mid-level planner then takes the output of the graph planner and assigns routes to each robot through the environment. Each group of robots moving together are then assigned roles in their coalition (i.e., leader or follower). This information is sent to each robot’s low-level planner to generate and execute dynamically feasible, collision-free paths.
Our technical approach is designed to address the high-level coordination problem presented in (2) and the local optimization problem described in Sec. III-B. To this end, we present a three-level planning hierarchy, where a top-level graph-based planner coordinates multi-robot teams, a mid-level A motion planner generates feasible paths between nodes, and low-level Model Path Predictive Integral Control (MPPI) [52] enables receding-horizon multi-robot motion planning. All three levels share underlying metric cost maps to maintain consistent planning across the hierarchy. Both A search and MPPI are surrogate approaches, selected for their widespread adoption, and could be substituted for other methods for path planning [53] and stochastic Nonlinear Model Predictive Control (NMPC) [54], respectively. This paper does not claim developments in these areas. Our key contributions are in the overall planning architecture and graph planning components.
While we believe our approach could apply more generally, in this paper, we focus on solutions to (2) and the local problem in Sec. III-B that solve the MCRP. We assume prior knowledge of the terrain and probability distributions for expected observer positions, as this information would have motivated the reconnaissance mission in this area. This information may be a coarse representation that evolves as we execute a plan or be provided with limited time before a scenario. In both cases, we require an approach for efficient planning to enable rapid response times to the given scenario and/or evolving information.
@@
Fig. 2 shows a flow diagram of our hierarchical planning architecture applied to the MCRP. First, we use our prior information to compute visibility scores, relative to the expected observer distributions, across an area of interest. Using the resulting visibility map, we segment the environment into regions of cover (or low visibility), which we use as nodes in a graph structure. We then calculate minimum visibility paths, using A search, between the regions to serve as edges in the graph. Between each node and edge pair, we compute overwatch scores based on the visibility to the edge from within the node. Together, this information forms a topological graph of the environment. We encode this graph in our MIP graph planner to compute an optimal plan for the multi-robot team. To form our mid-level plan, we assign each robot a route, which corresponds to a A path, and a role in the coalitions of robots that form (i.e., leader or follower). Each robot then controls to its path using MPPI in our low-level planner, performs overwatch, and/or moves in formation with their coalition. The following sections outline the algorithms that comprise each block of our overall hierarchical planner. Our core contribution is the graph planning algorithm in Sec. V. We introduce the construction of graphs for the MCRP in Sec. VI. Finally, we present the technical detail of our mid-level and low-level planners in Sections VII and VIII, respectively.
V Graph Planner
To generate an optimal plan for a team of robots coordinating on a graph, we developed a MIP approach to solve an instance of the problem presented in (2). In this section, we first define our key formulation innovations in Sec. V-A. We then detail how we utilize each innovation to formulate the specific cost function and constraints for the MCRP in Sections V-B and V-C, respectively. Ultimately, our derivation results in a statement of our optimization problem in Sec. V-D and a discussion of our formulation considerations in Sec. V-E.
V-A Key Formulation Innovations
A key contribution of this work is our computationally-efficient MILP optimization problem. We apply three core innovations to preserve linearity in our MILP formulation and significantly reduce the computation time to find an optimal solution to the problem: our integer decision variable formulation, utilization of indicator variables, and utilization of methods for formulating convex cost functions.
V-A1 Integer Decision Variable Formulation
Common formulations for multi-agent problems consider all possible states of robots at locations in the graph [15, 18]. For $n_{L}$ locations, this is $n_{L}^{n_{A}}$ possible states. In a MIP formulation, enumerating these states as binary variables would be computationally intractable. Another common formulation is to encode each robot’s location at each time step [25, 55]. This enumerates as $n_{T}n_{L}n_{A}$ total variables which is also highly computationally intensive for large numbers of agents and locations.
We drastically reduce the size of the decision space by instead considering the number of robots at each location at each time step with $n_{T}n_{L}$ integer variables. We then present methods for analyzing interactions between robots (e.g., sharing edges or providing overwatch) to maintain the same utility of the larger decision space, while eliminating the exponential growth. This formulation obfuscates the paths of each individual robot in the MILP problem, though we detail in Sec. VII-A how they can be assigned afterwards. Additionally, in the formulation of the optimization problem, we ensure that the paths follow the topology of the graph through sequential flow constraints considering the movement of all robots.
With this formulation, we embed the multi-robot interactions directly in the edge weights of our DTG. In particular, we consider benefits from moving in formations and as a team to directly scale with the number of robots on an edge; thus our choice of decision variable greatly simplifies these calculations. For interactions through overwatch opportunities, we define costs and constraints for each overwatch opportunity that depend on the number of robots at the overwatch nodes and related edges to reduce the cost of traversing those edges. Fully defining each of these possible interactions when instead enumerating variables for each robot would significantly increase the problem’s complexity. Our choice of decision variables is pivotal to keeping the problem computationally tractable in a complex decision space.
V-A2 Indicator Variables
In some cases, adding a variable for whether a condition is met aids in simplifying an expression or ensuring linearity. For this purpose, we use indicator variables [56], which are linked to another variable in the problem. The following proposition defines our form for indicator variables:
Proposition 1.
For $h\in[0,h^{\ulcorner}]$, an indicator variable $\alpha\in\{0,1\}$ can be expressed, with any constant $a\geq 0$, as follows:
$\alpha=\begin{cases}1,~~h>0\\ 0,~~h=0\end{cases}~~~\Leftrightarrow~~~~~\min a\alpha~\text{s.t.}~\alpha\geq\frac{h}{h^{\ulcorner}}.$ (3)
We can then use the newly created binary variable $\alpha$ in our cost functions and constraints. As long as $\alpha$ is used in contexts that will be minimized, the value of $\alpha$ will be strict to the piecewise expression in (3). This is due to the formulation; $\alpha$ can only be 0 or $1$, and the constraint follows that $\frac{h}{h^{\ulcorner}}>0$ for $h>0$ and $\frac{h}{h^{\ulcorner}}=0$ for $h=0$. For $\alpha$ to be minimized means that the associated cost function has the lowest value when $\alpha=0$ (compared to $\alpha=1$). Indicator variables of this form can aid in expressing complex equations more efficiently in our optimization problem.
V-A3 Convex Cost Functions
We aim to express our costs and constraints convexly, since convexity can significantly reduce the computation time of an optimization program. Since we are using MIP, we consider convexity when our integrality constraints are relaxed. For brevity, references to convexity of a function in this paper will assume this relaxation.
To produce a convex and positively homogeneous function, we can form a perspective function, as originally presented in [57]. In particular, we utilize the following result from [58]:
Proposition 2.
$\tilde{f}(x,y)=yf\Big(\frac{x}{y}\Big),\quad x\in\mathbb{R}^{n},~y\in\mathbb{R}_{>0}.$ (4)
In this work, we formulate piecewise cost functions based on the different regimes of operating alone or with a team. Using perspective functions, we can derive convex piecewise cost functions, which we will show in Sec. V-B1. To incorporate our piecewise functions linearly in our problem formulation, we can express these functions through a linear cost term with linear constraints using Proposition 3, which is commonly referred to as the epigraph form of the problem [59].
Proposition 3.
If $f(x)$ is a convex piecewise linear function with constraints $c_{1}(x),\dots,c_{n}(x)$, then the following are equivalent for some auxiliary variable $\breve{f}$:
$\min f(x)\quad\Leftrightarrow\quad\min\breve{f}~\text{s.t.}~\breve{f}\geq c_{1}(x),\dots,\breve{f}\geq c_{n}(x).$ (5)
V-B MIP Cost Function
We define the parameters that we use to formulate our MIP problem in Table I and the decision variables in Table II. Furthermore, we specify the constraints for the lower bounds (LB) and upper bounds (UB) of our decision variables in Table II. Using the information embedded in our DTG and the innovations described above, we formulate a cost function specific to the MCRP which minimizes the risk of traversing and total time to reach a set of target locations. The dynamic edge cost in our DTG, $w_{e,t}$ in Def. III.1, is a function of two cost functions we will define in the following section: (i) a cost of traversing $C_{\bar{W}_{e,t}}$, which embeds a fixed cost to traverse an edge, $\bar{w}_{e}$, and teaming considerations, and (ii) a cost reduction from overwatch $C_{\Omega_{{o},t}}$. In particular, $w_{e,t}=C_{\bar{W}_{e,t}}+C_{\Omega_{{o},t}}$.
(b) Cost of overwatch opportunity o{o} at time $t$ versus number of overwatch robots
Figure 3: Piecewise Linear Cost Functions
V-B1 Cost of Traversing
We formulate a cost of traversing with three main components: (a) edge cost, (b) vulnerability cost and (c) teaming rewards. Using the parameters from Table I, our base edge cost is the positive valued $\bar{w}_{e}$ for a team of robots to traverse edge $e$. For particularly vulnerable edges, it is desired for more agents to move together across the edge to enable moving in a formation for greater awareness. Thus, for each edge, to encode vulnerability, a minimum desired number of robots $a_{e}$ to traverse the edge is specified and an additional positive cost $m_{e}$ is incurred for each robot until the minimum is met. Finally, we consider an incentive for further teaming through positive rewards $r_{e}$ for additional robots moving together over the minimum desired. These components comprise the piecewise linear cost of traversing an edge at a particular time, $C_{\bar{W}_{e,t}}$, which can be expressed as a function of the number of robots on the edge at that time, $p_{e,t}$. This cost function is depicted in Fig. 3(a).
$C_{\bar{W}_{e,t}}=\begin{cases}0,&p_{e,t}=0\\ \bar{w}_{e}+m_{e}(a_{e}-p_{e,t}),&0<p_{e,t}\leq a_{e}\\ \bar{w}_{e}-r_{e}(p_{e,t}-a_{e}),&a_{e}<p_{e,t}\leq n_{A}.\end{cases}$ (6)